# Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models

### v2

@@ Abstract
Hallucinations present a significant challenge for large language models (LLMs). The utilization of parametric knowledge in generating factual content is constrained by the limited knowledge of LLMs, potentially resulting in internal hallucinations. While incorporating external information can help fill knowledge gaps, it also introduces the risk of irrelevant information, thereby increasing the likelihood of external hallucinations. To balance the use of parametric knowledge within LLMs and external information, in this study, we present Rowen, a novel framework that enhances LLMs with an adaptive retrieval augmentation process tailored to address hallucinated outputs. Rowen introduces a consistency-based hallucination detection module, which assesses the model’s uncertainty regarding the input query by evaluating the semantic inconsistencies in various responses generated across different languages or models. When high uncertainties in the responses are detected, Rowen activates the retrieval of external information to rectify the model outputs. Through comprehensive empirical experiments, we demonstrate that Rowen surpasses the current state-of-the-art in both detecting and mitigating hallucinated content within the outputs of LLMs111Our code is available at https://github.com/dhx20150812/Rowen..
Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models
Hanxing Ding1,2 Liang Pang1††thanks: Corresponding author Zihao Wei1,2 Huawei Shen1,2 Xueqi Cheng1,2 1Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
1 Introduction
@@
In recent years, large language models (LLMs) have demonstrated impressive abilities in natural language understanding Hendrycks et al. (2021); Huang et al. (2023c), generation Touvron et al. (2023); Taori et al. (2023), and reasoning Zhang et al. (2023e); Wang et al. (2023a); Chu et al. (2023). Despite their successes, it has been widely observed that even state-of-the-art LLMs often generate factually incorrect or nonsensical outputs, referred to as hallucinations Ji et al. (2023a); Zhang et al. (2023d, b). These unreliable outputs pose significant risks in practical deployments of LLMs.
@@
Efforts to enhance the factual accuracy of LLM outputs have been substantial. These studies often utilize LLMs’ extensive parametric knowledge and advanced logical reasoning capabilities. They employ approaches like self-reflection Wang et al. (2023b); Ji et al. (2023b); Madaan et al. (2023); Dhuliawala et al. (2023) or collaborative refinements involving interactions among multiple models Cohen et al. (2023); Du et al. (2023), aiming to enhance logical coherence in refined content. Despite their effectiveness, these self-improvement methods may be limited by LLMs’ knowledge boundaries Ren et al. (2023); Li et al. (2023a) or may not fully exploit parametric knowledge Huang et al. (2023a), leading to what we term internal hallucination, illustrated in Figure 1.
@@
Alongside these self-improvement strategies, Retrieval-Augmented Generation (RAG) Chern et al. (2023) serves as a complementary method to overcome knowledge limitations. RAG employs a retrieve-then-read pipeline Karpukhin et al. (2020); Lewis et al. (2020), integrating relevant documents from external knowledge sources into the LLMs’ generation process Huang et al. (2023a); Ye et al. (2023); Varshney et al. (2023); Chern et al. (2023); Li et al. (2023b). However, as depicted in Figure 1, RAG methods are susceptible to external hallucination when irrelevant evidence is incorporated, potentially leading to cumulative errors and compromising output accuracy Li et al. (2023a); Shi et al. (2023).
@@
Drawing inspiration from the latest neuroscience research Poskanzer and Aly (2023), which reveals how the human brain dynamically switches between internal thoughts and external sensations, we introduce a novel method, termed Rowen (Retrieve only when it needs). Rowen involves an innovative consistency-based uncertainty estimation module, which perturbs semantically equivalent questions and then evaluates the semantic inconsistency of various responses across diverse languages and models when subjected to these perturbed queries. To mitigate internal hallucinations, we trigger a retrieval process to fetch relevant information when Rowen detects inconsistencies in LLMs’ responses, indicating internal reasoning failures. This helps LLMs refine their reasoning chains and rectify potential hallucinations. To reduce external hallucinations, Rowen minimizes the risk of incorporating erroneous information by optimizing retrieval phases. If the perturbed answers convey consistent content, suggesting that LLMs are capable of generating the correct answer themselves, we directly adopt the original answer produced by internal reasoning. This method integrates parametric knowledge within LLMs with retrieved sources, ensuring a balanced integration of internal reasoning and external evidence to effectively mitigate hallucinations.
@@
We evaluate the effectiveness of Rowen on the TruthfulQA dataset Lin et al. (2022) and the StrategyQA dataset Geva et al. (2021). Remarkably, our approach excels on the TruthfulQA dataset, yielding an impressive GPT-Judge score of 59.34%, marking a substantial improvement over the state-of-the-art (SOTA) baseline by a significant margin (+16.74%). Similarly, on the StrategyQA dataset, our approach achieves an accuracy of 75.60%, surpassing existing self-improvement and RAG-based baselines with notable superiority. These results comprehensively underscore the powerful capability of Rowen in mitigating hallucinated outputs of LLMs. Furthermore, our adaptive retrieval strategy significantly reduces unnecessary retrievals, thereby enhancing the efficiency of RAG systems.
2 Related Works
@@
Mitigating hallucinations in the inference time could be a cost-effective and controllable way. A line of research harnesses the extensive parametric knowledge and robust logical reasoning capabilities of LLMs to ensure logical consistency either through self-reflection within a single model Wang et al. (2023b); Ji et al. (2023b); Xu et al. (2023); Madaan et al. (2023); Dhuliawala et al. (2023) or through collaborative refinements or debates involving multiple models Cohen et al. (2023); Du et al. (2023). Despite their strengths, LLMs are sometimes constrained by their knowledge boundaries and the complexity of the reasoning chain, resulting in occasional inaccuracies Ren et al. (2023); Li et al. (2023a); Mallen et al. (2023) termed internal hallucination. To address this knowledge gap, retrieval-augmented generation methods leverage external knowledge as supplementary evidence to aid LLMs in providing accurate responses Huang et al. (2023a); Ye et al. (2023); Niu et al. (2012); Xu et al. (2024); Varshney et al. (2023); Chern et al. (2023); Li et al. (2023b). However, these approaches, while effective, occasionally encounter the challenge of error accumulation, where irrelevant evidence may seep into the generation process, leading to incorrect responses Li et al. (2023a); Shi et al. (2023), a phenomenon referred to as external hallucination. Our work only performs retrieval augmentation when hallucinations are detected, thereby maximizing the utilization of both the parametric knowledge and externally retrieved information.
There are also some adaptive retrieval methods that assess the difficulty of questions or the confidence in responses to decide whether to retrieve documents Jiang et al. (2023); Mallen et al. (2023); Asai et al. (2023); Jeong et al. (2024). Due to space limitations, we discuss these works in Appendix C.
3 Methodology
@@
3.1 Stage 1: Generating Initial Answer
To maximize the exploitation of the parametric knowledge in LLMs, we initially employ their Chain-of-Thought (CoT) reasoning to generate a preliminary response. This process involves: critically assessing the validity of the information in the input query and prioritizing accuracy and fact-checking for answer elaboration, detailed in Appendix F Table 9. After generating the CoT thought , we ask the model to provide a concise answer for the input query . The answer will serve as the final response to the input query if our detection module ensures it is free from hallucinations.
3.2 Stage 2: Deciding Whether to Retrieve
To decide when to retrieve, we leverage model uncertainty, which refers to the confidence level of model outputs and serves as a crucial indicator for deciding when to trust LLMs Zhang et al. (2023d). Unfortunately, current consistency-based methods fail when LLMs provide consistent yet incorrect answers across different perturbations Zhang et al. (2023a). This issue may arise because these methods focus exclusively on semantic coherence within a single language or model.
@@ 3.2 Stage 2: Deciding Whether to Retrieve
3.2.1 Diversifying Question Verbalizations
To facilitate subsequent consistency-based hallucination detection, we begin by leveraging advancements in LLM prompting to generate semantically equivalent perturbations. Initially, we start with an input and instruct the model to provide a set of semantically equivalent questions . We use the prompt: "For the question [ORIGINAL QUESTION], please provide k semantically equivalent questions" with a high decoding temperature to generate diverse perturbed expressions.
@@ 3.2.1 Diversifying Question Verbalizations
3.2.2 Cross-Language Consistency-based Detection
Recent finding reveals that multilingual models do not always learn knowledge that is fully aligned across different languages Gao et al. (2024). Building on this insight, we propose that this inconsistency can be strategically used as an indicator for the necessity of information retrieval. When language models provide inconsistent answers to the same query in different languages, it signals a lack of certainty about the relevant knowledge. In such cases, retrieving external knowledge is crucial for accurate responses.
To achieve cross-language detection, we first incorporate a cross-language consistency check (Rowen-CL) to evaluate the semantic consistency of responses to the same question across different languages. We introduce language-level perturbations by asking the model to translate the source-language questions into corresponding paraphrased questions in the target language. The model is then instructed to generate corresponding answers to each question in the target language.
To captures language-level cross-checking uncertainties, we utilize the generated questions and answers from all previous stages to calculate a numerical consistency score to calculate a numerical consistency score. Let denote a semantic equivalence checking operator that takes two QA pairs as inputs and returns "True" if they are semantically equivalent, and "False" otherwise. We then map the response to a numerical semantic equivalent score: {"True" 1.0, "False" 0.0}. We leverage the model and utilize the prompt provided in Appendix F Table 10 to implement the cross-language checking operator and calculate the cross-language consistency score as:
@@ 3.2.2 Cross-Language Consistency-based Detection
3.2.3 Cross-Model Consistency-based Detection
Besides language-level cross-checking, we also introduce a cross-model detection module (Rowen-CM) to evaluate the semantic consistency of responses to the same question across different models. We adopt an additional verifier LM for model-level cross-checking and instruct the verifier LM to provide answers for each semantically equivalent question:
3.2.4 Hybrid Consistency-based Detection
@@
3.3 Stage 3: Retrieval Augmented Generation
@@
To help the LM correct errors, we search for supporting evidence from external sources like online webpages. We first ask the model to generate search queries for each paraphrased question in , following the prompt in Appendix F Table 12. These queries are input into the online search engine to retrieve relevant knowledge, denoted as , used for correcting factual errors in .
@@
Appendix F Table 13 shows the prompt for repairing the original answer. The corrected answer serves as the final response to question .
4 Experimental Setup
4.1 Datasets and Evaluation Metrics
TruthfulQA
We use the TruthfulQA dataset Lin et al. (2022) to evaluate the ability of LLMs in generating truthful responses Zhang et al. (2023c); Kadavath et al. (2022). In our study, we focus on the generation task in TruthfulQA. Therefore, to evaluate the factuality of responses from LLMs, we calculate the GPT-Judge score, obtained by fine-tuning the babbage-002 model using the original fine-tuning data from their official repository222https://github.com/sylinrl/TruthfulQA. We also report the BLEU and Rouge-L scores to evaluate the lexical overlap between generated responses and ground-truth references.
StrategyQA
The StrategyQA dataset Geva et al. (2021) comprises crowdsourced yes / no questions that require multi-step reasoning for accurate answers. We follow previous work Jiang et al. (2023) to randomly sample 500 examples due to the cost consideration of running experiments. We also follow the settings of Wei et al. (2022) to generate both the reasoning process as well as the final answer. We present the exact-match accuracy of the generated yes / no answers compared to the gold-standard answers.
@@ 4.1 Datasets and Evaluation Metrics
| CoVe Dhuliawala et al. (2023) | 48.01 | 12.81 | 26.52 | 61.40 |
| Multi-agent Debate Du et al. (2023) | 50.83 | 3.94 | 21.05 | 65.73 |
| Self-Reflection Ji et al. (2023b) | 42.99 | 3.86 | 18.18 | 62.40 |
@@
| Factool Chern et al. (2023) | 34.50 | 1.34 | 12.22 | 67.20 |
| Detect-and-Mitigate Varshney et al. (2023) | 49.98 | 3.17 | 18.59 | 56.94 |
@@
| FLARE Jiang et al. (2023) | 45.04 | 11.59 | 26.83 | 61.19 |
| Adaptive-Retrieval Mallen et al. (2023) | 45.55 | 8.87 | 26.75 | 62.50 |
| Self-RAG Asai et al. (2023) | 40.36 | 4.36 | 21.28 | 58.40 |
| Adaptive-RAG Jeong et al. (2024) | 46.02 | 10.29 | 26.24 | 68.50 |
| LUQ Zhang et al. (2024) | 55.08 | 5.79 | 21.44 | 71.00 |
@@
4.2 Baseline Methods

### v3

# Rowen: Adaptive Retrieval-Augmented Generation for Hallucination Mitigation in LLMs

@@ Abstract.
Hallucinations present a significant challenge for large language models (LLMs). The utilization of parametric knowledge in generating factual content is constrained by the limited knowledge of LLMs, potentially resulting in internal hallucinations. While incorporating external information can help fill knowledge gaps, it also introduces the risk of irrelevant information, thereby increasing the likelihood of external hallucinations. To balance the use of parametric knowledge within LLMs and external information, in this study, we present Rowen, a novel framework that enhances LLMs with an adaptive retrieval augmentation process tailored to address hallucinated outputs. Rowen introduces a consistency-based hallucination detection module, which assesses the model’s uncertainty regarding the input query by evaluating the semantic inconsistencies in various responses generated across different languages or models. When high uncertainties in the responses are detected, Rowen activates the retrieval of external information to rectify the model outputs. Through comprehensive empirical experiments, we demonstrate that Rowen surpasses the current state-of-the-art in both detecting and mitigating hallucinated content within the outputs of LLMs111Our Code: https://github.com/dhx20150812/Rowen..
1. Introduction
@@
In recent years, large language models (LLMs) have demonstrated impressive abilities in natural language understanding (Hendrycks et al., 2021; Huang et al., 2023a), generation (Touvron et al., 2023; Taori et al., 2023), and reasoning (Zhang et al., 2023e; Wang et al., 2023a; Chu et al., 2023). Despite their successes, it has been widely observed that even state-of-the-art LLMs often generate factually incorrect or nonsensical outputs, referred to as hallucinations (Ji et al., 2023a; Zhang et al., 2023b, d). These unreliable outputs pose significant risks in practical deployments of LLMs.
@@
Efforts to enhance the factual accuracy of LLM outputs have been substantial. These studies often utilize LLMs’ extensive parametric knowledge and advanced logical reasoning capabilities. They employ approaches like self-reflection (Wang et al., 2023b; Ji et al., 2023b; Madaan et al., 2023; Dhuliawala et al., 2023) or collaborative refinements involving interactions among multiple models (Cohen et al., 2023; Du et al., 2023), aiming to enhance logical coherence in refined content. Despite their effectiveness, these self-improvement methods may be limited by LLMs’ knowledge boundaries (Ren et al., 2023; Li et al., 2023b) or may not fully exploit parametric knowledge (Huang et al., 2023c), leading to what we term internal hallucination, illustrated in Figure 1.
@@
Alongside these self-improvement strategies, Retrieval-Augmented Generation (RAG) (Chern et al., 2023) serves as a complementary method to overcome knowledge limitations. RAG employs a retrieve-then-read pipeline (Karpukhin et al., 2020; Lewis et al., 2020), integrating relevant documents from external knowledge sources into the LLMs’ generation process (Huang et al., 2023c; Ye et al., 2023; Varshney et al., 2023; Chern et al., 2023; Li et al., 2023a). However, as depicted in Figure 1, RAG methods are susceptible to external hallucination when irrelevant evidence is incorporated, potentially leading to cumulative errors and compromising output accuracy (Li et al., 2023b; Shi et al., 2023).
@@
Drawing inspiration from the latest neuroscience research (Poskanzer and Aly, 2023), which reveals how the human brain dynamically switches between internal thoughts and external sensations, we introduce a novel method, termed Rowen (Retrieve only when it needs). Rowen involves an innovative consistency-based uncertainty estimation module, which perturbs semantically equivalent questions and then evaluates the semantic inconsistency of various responses across diverse languages and models when subjected to these perturbed queries. To mitigate internal hallucinations, we trigger a retrieval process to fetch relevant information when Rowen detects inconsistencies in LLMs’ responses, indicating internal reasoning failures. This helps LLMs refine their reasoning chains and rectify potential hallucinations. To reduce external hallucinations, Rowen minimizes the risk of incorporating erroneous information by optimizing retrieval phases. If the perturbed answers convey consistent content, suggesting that LLMs are capable of generating the correct answer themselves, we directly adopt the original answer produced by internal reasoning. This method integrates parametric knowledge within LLMs with retrieved sources, ensuring a balanced integration of internal reasoning and external evidence to effectively mitigate hallucinations.
@@
We evaluate the effectiveness of Rowen on the TruthfulQA dataset (Lin et al., 2022) and the StrategyQA dataset (Geva et al., 2021). Remarkably, our approach excels on the TruthfulQA dataset, yielding an impressive GPT-Judge score of 59.34%, marking a substantial improvement over the state-of-the-art (SOTA) baseline by a significant margin (+16.74%). Similarly, on the StrategyQA dataset, our approach achieves an accuracy of 75.60%, surpassing existing self-improvement and RAG-based baselines with notable superiority. These results underscore the powerful capability of Rowen in mitigating hallucinated outputs of LLMs. Furthermore, our adaptive retrieval strategy significantly reduces unnecessary retrievals, thereby enhancing the efficiency of RAG systems.
2. Related Works
In this section, we discuss recent works on hallucination detection and mitigation, focusing on uncertainty estimation methods for detection and post-hoc correction for mitigation.
2.1. Exploring Uncertainty for Hallucination Detection
Uncertainty refers to the confidence level of the model outputs, and it serves an important indicator for identifying and eliminating hallucinations, so it can assist users in determining when to trust LLMs. In general, methods for exploring uncertainty for hallucination detection can be categorized into three types: (1) Logit-based estimation relies on accessing the model’s logits to calculate token-level probabilities or entropy, which are used to measure uncertainty (Manakul et al., 2023; Huang et al., 2023b; Varshney et al., 2023). However, this approach can pose challenges for black-box closed-source models. (2) Verbalized-based estimation involves prompting language models to express their uncertainty using specific prompts (Xiong et al., 2023; Tian et al., 2023; Agrawal et al., 2023). However, these methods tend to display a high degree of overconfidence when expressing their confidence (Xiong et al., 2023; Tao et al., 2024). (3) To overcome these limitations, consistency-based estimations are adopted to measure the consistency score among multiple responses provided by the model for a given question (Manakul et al., 2023; Xiong et al., 2023; Wang et al., 2023b; Zhao et al., 2023). The underlying assumption suggests that when language models struggle with indecision and fabricate facts, they tend to provide logically inconsistent responses to identical questions. In this work, we propose that cross-language and cross-model consistency can offer highly sensitive signals for identifying hallucinations. Therefore, we utilize cross-language and cross-model detection modules that cross-check answers to the same question across different languages or models. This cross-checking paradigm serves as a powerful mechanism to identify hallucinations in LLMs.
2.2. Post-hoc Correction for Hallucination Mitigation
@@
Mitigating hallucinations in the inference time could be a cost-effective and controllable way. A line of research harnesses the extensive parametric knowledge and robust logical reasoning capabilities of LLMs to ensure logical consistency either through self-reflection within a single model (Wang et al., 2023b; Ji et al., 2023b; Xu et al., 2023; Madaan et al., 2023; Dhuliawala et al., 2023) or through collaborative refinements or debates involving multiple models (Cohen et al., 2023; Du et al., 2023). Despite their strengths, LLMs are sometimes constrained by their knowledge boundaries and the complexity of the reasoning chain, resulting in occasional inaccuracies (Ren et al., 2023; Li et al., 2023b; Mallen et al., 2023) termed internal hallucination. To address this knowledge gap, retrieval-augmented generation methods leverage external knowledge as supplementary evidence to aid LLMs in providing accurate responses (Huang et al., 2023c; Ye et al., 2023; Niu et al., 2012; Xu et al., 2024; Varshney et al., 2023; Chern et al., 2023; Li et al., 2023a). However, these approaches, while effective, occasionally encounter the challenge of error accumulation, where irrelevant evidence may seep into the generation process, leading to incorrect responses (Li et al., 2023b; Shi et al., 2023), a phenomenon referred to as external hallucination. Our work only performs retrieval augmentation when hallucinations are detected, thereby maximizing the utilization of both the parametric knowledge and externally retrieved information.
3. Methodology
@@
3.1. Stage 1: Generating Initial Answer
To maximize the exploitation of the parametric knowledge in LLMs, we initially employ their Chain-of-Thought (CoT) reasoning to generate a preliminary response. This process involves: critically assessing the validity of the information in the input query and prioritizing accuracy and fact-checking for answer elaboration, detailed in Table 1. After generating the CoT thought , we ask to provide a concise answer for the input query . Our aim is to ensure high-quality responses are generated from the outset. The answer is finalized as the ultimate response to the input query after our detection module ensures it is free from hallucinations.
3.2. Stage 2: Deciding Whether to Retrieve
To decide when to retrieve, we leverage model uncertainty, which refers to the confidence level of model outputs and serves as a crucial indicator for deciding when to trust LLMs (Zhang et al., 2023b). Unfortunately, current consistency-based methods fail when LLMs provide consistent yet incorrect answers across different perturbations (Zhang et al., 2023c). This issue may arise because these methods focus exclusively on semantic coherence within a single language or model.
@@ 3.2. Stage 2: Deciding Whether to Retrieve
3.2.1. Cross-Language / Model Perturbations
To facilitate subsequent consistency-based hallucination detection, we begin by leveraging advancements in LLM prompting to generate semantically equivalent perturbations. Initially, we start with an input and instruct the model to provide a set of semantically equivalent questions . We use the prompt: ”For the question [ORIGINAL QUESTION], please provide k semantically equivalent questions” with a high decoding temperature to generate diverse perturbed expressions.
@@ 3.2.1. Cross-Language / Model Perturbations
Cross-Language Detection
To capture language-level uncertainty, we first incorporate a cross-language consistency check (Rowen-CL) to evaluate the semantic consistency of responses to the same question across different languages. To achieve this, we introduce language-level perturbations by asking the model to translate the source-language questions into corresponding paraphrased questions in the target language. The model is then instructed to generate corresponding answers to each question in the target language.
Cross-Model Detection
Besides language-level cross-checking, we also introduce a cross-model detection module (Rowen-CM) to evaluate the semantic consistency of responses to the same question across different models. We adopt an additional verifier LM for model-level cross-checking and instruct the verifier LM to provide answers for each source question:
3.2.2. Consistency Score Calculation
In this step, we utilize the generated questions and answers from all previous stages to calculate a numerical consistency score that captures language-level and model-level cross-checking uncertainties.
Cross-Language Consistency Score
Let denote a semantic equivalence checking operator that takes two QA pairs as inputs and returns ”True” if they are semantically equivalent, and ”False” otherwise. We then map the response to a numerical semantic equivalent score: {"True" 1.0, "False" 0.0}. In our implementation, we leverage the model and utilize the prompt provided in Table 2 to implement the cross-language checking operator and calculate the cross-language consistency score as:
@@ 3.2.2. Consistency Score Calculation
Cross-Model Consistency Score
Similar to the cross-language consistency score calculation, we use the prompt in Table 3 to implement the checking operator to calculate cross-model consistency score:
where denote the QA pairs generated by verifier model .
Hybrid Consistency Score
@@
3.3. Stage 3: Retrieval Augmented Generation
@@
To help the LM correct errors, we search for supporting evidence from external sources like online webpages. We first ask the model to generate search queries for each paraphrased question in . These queries are input into the online search engine to retrieve relevant knowledge, denoted as , used for correcting factual errors in .
@@
The corrected answer serves as the final response to question .
@@
| CoVe (Dhuliawala et al., 2023) | 48.01 | 12.81 | 26.52 | 61.40 |
| Multi-agent Debate (Du et al., 2023) | 50.83 | 3.94 | 21.05 | 65.73 |
| Self-Reflection (Ji et al., 2023b) | 42.99 | 3.86 | 18.18 | 62.40 |
@@
| Factool (Chern et al., 2023) | 34.50 | 1.34 | 12.22 | 67.20 |
| Detect-and-Mitigate (Varshney et al., 2023) | 49.98 | 3.17 | 18.59 | 56.94 |
@@
| FLARE (Jiang et al., 2023) | 45.04 | 11.59 | 26.83 | 61.19 |
| Adaptive-Retrieval (Mallen et al., 2023) | 45.55 | 8.87 | 26.75 | 62.50 |
| Self-RAG (Asai et al., 2023) | 40.36 | 4.36 | 21.28 | 58.40 |
| Adaptive-RAG (Jeong et al., 2024) | 46.02 | 10.29 | 26.24 | 68.50 |
| LUQ (Zhang et al., 2024) | 55.08 | 5.79 | 21.44 | 71.00 |
@@
4. Experimental Setup
In this section, we outline the experimental setup of Rowen, detailing the datasets utilized, the evaluation metrics employed, the baseline methods considered, and the implementation specifics.
4.1. Datasets and Evaluation Metrics
We evaluate the hallucination mitigation performance on the following datasets:
TruthfulQA
We use the TruthfulQA dataset (Lin et al., 2022) to evaluate the ability of LLMs in generating truthful responses (Zhang et al., 2023a; Kadavath et al., 2022). In our study, we focus on the generation task in TruthfulQA. Therefore, to evaluate the factuality of responses from LLMs, we calculate the GPT-Judge score, obtained by fine-tuning the babbage-002 model using the original fine-tuning data from their official repository222https://github.com/sylinrl/TruthfulQA. We also report the BLEU and Rouge-L scores to evaluate the lexical overlap between generated responses and ground-truth references.
StrategyQA
The StrategyQA dataset (Geva et al., 2021) comprises crowdsourced yes / no questions that require multi-step reasoning for accurate answers. We follow previous work (Jiang et al., 2023) to randomly sample 500 examples due to the cost consideration of running experiments. We also follow the settings of Wei et al. (2022) to generate both the reasoning process as well as the final answer. We present the exact-match accuracy of the generated yes / no answers compared to the gold-standard answers.
4.2. Baseline Methods